OpenStudy (anonymous):

simplify 24x^3 over 8x^-2

OpenStudy (anonymous):

You should be able to work this out based on the explanation i gave you earlier, or do you just want someone to give you the answer.

OpenStudy (anonymous):

i didnt get u explanation earlyer so...ya..

OpenStudy (anonymous):

Yes I did, I typed out a whole mess of explanations, if you bothered to look.

OpenStudy (anonymous):

i looked at it for 5 min i didnt understand it..

OpenStudy (anonymous):

well then don't say i didn't give an explanation say I didn't get it. What was the first step I told you to take?

OpenStudy (anonymous):

make it in2 tiny parts

OpenStudy (anonymous):

right, factor it. Break it down into it's smallest numbers. So what do you get when you break it down like that?

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

\[24x^3/8x^{-2}\] \[3x^3 \times x^2 = 3x^{2 + 3} = 3x^5\] Because: \[x^m \times x^n = x^{m + n}\] And, \[1/x^{-n} = x^n\]